OpenStudy (faiqraees):

Let \(I_n = \int_{0}^{1} \frac{dx}{(1+x^2)^n} \) Prove that for every positive integer n. \(2nI_{n+1} = 2^{-n}+(2n-1)I_n \)

OpenStudy (faiqraees):

@ganeshie8

OpenStudy (faiqraees):

@IrishBoy123

OpenStudy (faiqraees):

I tried multiplying the integral by 1 and then using integration by parts bu ended with \[ I_n = 2^{-n} +2n \int_{0}^{1} \frac{x^2}{(1+x^2)^{n+1}} dx\] I then tried making the numerator 1+x^2-x^2 and ended up with \[ I_n = I_{n-1} + \int_{0}^{1} \frac{x^2}{(1+x^2)^{n}} dx\] What should I do?

ganeshie8 (ganeshie8):

\[ I_n = 2^{-n} +2n \int_{0}^{1} \frac{x^2}{(1+x^2)^{n+1}} dx\] Looks good till above step

OpenStudy (faiqraees):

yeah but there is a x^2 what should I do with it?

ganeshie8 (ganeshie8):

Next you may replace x^2 by 1+x^2-1

OpenStudy (faiqraees):

ohhhhhhh

OpenStudy (faiqraees):

gonna kill myself

ganeshie8 (ganeshie8):

happens..

OpenStudy (faiqraees):

thank you very much

ganeshie8 (ganeshie8):

Np